OpenStudy (anonymous):

For his phone service, Ahmad pays a monthly fee of $20, and he pays an additional $0.07 per minute of use. The least he has been charged in a month is $97.84. What are the possible numbers of minutes he has used his phone in a month? Use m for the number of minutes, and solve your inequality for m.

OpenStudy (anonymous):

@icecreamman

OpenStudy (anonymous):

97.84 - 20 = 77.84 divided by 0.07 =1,100?

OpenStudy (anonymous):

wait where did the 1100 come from?

OpenStudy (anonymous):

first set up your inequality

OpenStudy (anonymous):

dont jump the gun here start iwth the inequalit

OpenStudy (anonymous):

i subtracted 20 from 97.84 = 77.84 because thats the mmounthly fee and then i divided 77.84 by 0.07 for the minutes used?

OpenStudy (anonymous):

how should is set it up?

OpenStudy (anonymous):

set up your inequaltiy i want to see what you started with

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

you are sort of hinted that you needed to start with an equation

OpenStudy (anonymous):

i thought that is how you solve these problems

OpenStudy (anonymous):

no you have to start iwth the inequaltiy. so i will give you the starting point

OpenStudy (omarbirjas):

Ok so we know he has a base payment is 20. and for every minute you pay .07 cents. lets put that into an epression and we get \[20+.07m\] let m represent the number of mins.

OpenStudy (anonymous):

\[20+0.07m<91.84\]

OpenStudy (omarbirjas):

No @icecreamman

OpenStudy (anonymous):

how should i solve this?

OpenStudy (omarbirjas):

its greater than or equal to 91.84

OpenStudy (anonymous):

no it has the <97.84 because at the end of the problem it says The least he has been charged in a month is $97.84.

OpenStudy (omarbirjas):

\[20+.07m \ge 91.84\]

OpenStudy (omarbirjas):

Yea it says LEAST not MOST...

OpenStudy (omarbirjas):

now all you have to do is, subtract 20 from both sides then devide both sides by .07

OpenStudy (anonymous):

no i swear its least not "most " the sign you have is greater than or equal to i think that least means less than or equal to omar

OpenStudy (anonymous):

so im sure that it is \[20+0.07m<97.84\]

OpenStudy (omarbirjas):

91.84 is the smallest amount he's been charged....hence making the other side of the inequality more than that. and it HAS to be equal to because he's been charged that much.

OpenStudy (anonymous):

ok so then the inequality sign will be 

OpenStudy (omarbirjas):

@iambatman HALP US, we haz a dis agreements.

OpenStudy (omarbirjas):



OpenStudy (anonymous):

so 97.84-20 = 77.84

OpenStudy (anonymous):

yes @RaeBeach9296

OpenStudy (omarbirjas):

no...

OpenStudy (omarbirjas):

icecreamman is trying to give you the wrong answer....

OpenStudy (anonymous):

now divide 77.84 by 0.07? = 1,112?

OpenStudy (omarbirjas):

@confluxepic

OpenStudy (anonymous):

Lol, it's greater than equal to, it's at least.

OpenStudy (anonymous):

@iambatman can you please explain?

OpenStudy (omarbirjas):

See ^-^

OpenStudy (anonymous):

now im confuzed

OpenStudy (omarbirjas):

any whooooo 20+.07m≥91.84 now subtract 20 from both sides and devide by .07, what do you get?

OpenStudy (anonymous):

1,112

OpenStudy (omarbirjas):

Mr.WolframAlpha speaks to me otherwise fellow lad...

OpenStudy (omarbirjas):

subtract 20 from both sides and you get .07m≥71.84

OpenStudy (omarbirjas):

\[.07m \ge 71.84\]

OpenStudy (omarbirjas):

Multiply both sides by a hundered to make it simpler and you get \[7m \ge 7184\]

OpenStudy (omarbirjas):

now devide both sides by seven and you get \[m \ge 1026.285714..............\]

OpenStudy (omarbirjas):

Tada